A proton travels at a uniform speed of 2.50 × 106 m/s toward the right of this page. It then enters a uniform magnetic field and experiences a maximum force of 4.95 × 10-8 N that points toward the top of the page. What is the magnitude of the magnetic field?What is the direction of the magnetic field?
Answer #1

The equation for the force is

F = Bq v

B =\frac{F}{qv} = \frac{4.98 \times 10^{-8}}{(1.6 \times 10^{-19})(2.5 \times 10^6)} = 1.245 \times 10^5 T

The force is upward, speed is to the right so the magnetic field is into the page as per RHR